# Asymptotic behavior of the ground state energy of a fermionic Fröhlich multipolaron in the strong coupling limit

### v2

@@ 1.3 Statement of main result: Strong-coupling limit
1.2 Remark.
We note here that, in contrast to [Wel15], we do not assume that
$E_{PT}^{(n,\alpha)}\leq E_{PT}^{(k,\alpha)}+E_{PT}^{(n-k,\alpha)},\quad\forall n\leq N,\quad\forall k<n.$ (1.24)
This is an important improvement as it is not clear when this assumption is satisfied unless the external electric and magnetic fields are constant or periodic.
1.3 Remark.
$E_{PT}^{(N,\alpha)}(\Psi_{\alpha})\,=\,\alpha^{2}E_{PT}^{(N,1)}(\Psi)\,.$ (1.25)
Consequently, the Pekar-Tomasevich energy $E_{PT}^{(N,\alpha)}=\alpha^{2}E_{PT}^{(N,1)}$ is, in fact, the leading order term in the RHS of (1.23) in Theorem 1.1, as long as $E_{PT}^{(N,1)}\neq 0$.
@@
where $x_{+}=\max(x,0)$, see Lemma C.1. The infimum in (1.27) is finite, since the electron-electron interaction grows at most with $N^{2}$, while the rest grows to linear order. To bound the terms, one employs Hardy’s inequality together with the diamagnetic inequality [LL96, Theorem 7.21], and the assumptions on $A$ and $V$.
@@
In the absence of external fields and in the physically relevant regime $\nu>2$, Griesemer and Møller [GMl10] established stability of the Pekar-Tomasevich functional in the sense that
@@
see also [FLST11]. This bound is used only in the auxiliary Lemma C.1, see also (Block-mode reduction)., and improves the quadratic growth, see Remark 1.4. As a consequence, one can improve the bound the $N$-dependence of the error in Theorem 1.1 to
@@
for some constant $C_{\alpha}$ independent of $N$, see [FLST11, Theorem 3]. In the regime $\nu<2$, this stability bound ceases to be true. In fact, by [GMl10], one has that
@@
As a consequence of Theorem 1.1, one can establish binding of the multipolaron in the case that $V=0$ and $A$ is linear, corresponding to a constant magnetic field, as along as as $\nu<\nu_{c}(A,N)$ for some threshold $\nu_{c}(A,N)>2$. More precisely, one of the authors and Griesemer [AG14] proved the existence of $\nu_{c}(A,N)>2$ s.t. for all $\nu<\nu_{c}(A,N)$
@@
Lewin [Lew11] proved an analogous result in the case $A=0$. As a consequence of this result and Theorem 1.1, one obtains for all $\alpha>0$ sufficiently large
@@
As in [GW13], one can prove that without rescaling $A\to A_{\alpha}$, $V\to V_{\alpha}$, $\lim_{\alpha\to\infty}\alpha^{-2}E^{(N,\alpha)}$ is independent of the external fields, under sufficient regularity of $A$ and $V$.
Sketch of the proof of Theorem 1.1
In this work, we follow the overall strategy in [Wel15] but, crucially, we take care of the fermionic statistics. To this end, we employ a localization method developed in [LL05]. We also refine the proof of [Wel15] in order to relax the assumption (1.24) on the external fields, to allow for a large class of external fields.
The paper is organized as follows: In Section 2, we start by proving that the electrons can be localized in clusters of well-separated balls of comparable radius with controllable error using a localization function found in [LL05] and a covering argument found in [Wel15]. As in [Wel15], we show that the interaction between clusters is negligible if the radius of the balls is chosen to be $R=N^{-\frac{11}{30}}\alpha^{-\frac{19}{23}}$, and $\alpha>0$ large.
Then, we invoke a phonon localization result from [Wel15], which generalizes the single electron case studied [LT97], as well as the bipolaron case studied in [FLST11] to many electrons. It allows us to reduce the phonon modes to those closest to the respective electronic clusters. To coarse-grain the energy of the individual clusters to the Pekar-Tomasevich energy, we repeat the arguments in [LT97]. They consist of first UV-truncating the Hamiltonian, Lemma 3.1 and partitioning the truncated Hamiltonian into phonon block modes, Prop. 3.2. Let us abbreviate the resulting reduced Hamiltonian $H_{\mathrm{block}}$, see (3.7) and (3.9). The crucial insight of [LT97] lies in realizing that one can then expand the polaron wave function in terms of products with phonon coherent states over the finitely many block modes involved in $H_{\mathrm{block}}$. This expansion allows us to repeat Pekar’s argument, see Lemma A.1, applied to $H_{\mathrm{block}}$, see Proposition 3.3. We sketch out the main idea, which is at the core of Lieb-Thomas’ analysis, after Lemma A.1 in the appendix.
One of our new insights is that when reducing to the Pekar-Tomasevich energy, one can preserve the electronic localization within balls. This, in turn, enables us to bound the sum over the energy of the individual, non-interacting clusters by the total Pekar-Tomasevich energy of the interacting electrons. This is why we can remove assumption (1.24) from [Wel15].
2 Localization of the Fröhlich multipolaron
In the following, we identify $\Psi\in\mathcal{H}_{N}$ with a vector $\iota(\Psi)\in L^{2}(\mathbb{R}^{3N};\mathcal{F})$ and denote by $\operatorname{supp}_{e}(\Psi):=\operatorname{supp}(\iota(\Psi))$ its electronic support. As usual, we denote by $B_{r}(x)\subseteq\mathbb{R}^{3}$ the open ball of radius $r$ centered at $x$.
Our goal is to localize an approximate minimizer of the Fröhlich Hamiltonian in a union of balls that far from each other. The constructed minimizer will be antisymmetric within clusters contained in distinct balls, but not among different balls. Due to the spatial separation of the balls, this will allow us a decomposition of the Hamiltonian into independent clusters.
As explained, we repeat a localization procedure taken from [LL05]. In particular, given $\chi$ a non-negative, $L^{2}$-normalized test function localized in $B_{R}(0)$, define for all $X=(x_{1},\ldots,x_{N})$, $Y=(y_{1},\ldots,y_{N})$
@@ 2 Localization of the Fröhlich multipolaron
Let $\Phi\in\mathcal{H}_{N}$ be normalized. With steps analogous to those in [LL05, Lemma 4.1], one can show that $W$ is well-defined and smooth, and that there exists some $Y_{0}\in\mathbb{R}^{3N}$ such that $W_{Y_{0}}\Phi\neq 0$, and
@@
Next, we use a covering lemma taken from [Wel15] to partition electrons into balls of possibly larger size that are well-separated.
2.1 Lemma (Covering lemma).
  1. (i)
    $B_{j}$ has radius $R_{j}=\frac{1}{2}(3n_{j}-1)R\quad\forall 1\leq j\leq m$,
  2. (ii)
    $\operatorname{dist}(B_{i},B_{j})\geq R$ for $i\neq j$,
  3. (iii)
    $\bigcup_{k\in C_{j}}B_{R}(y_{k})\subseteq B_{j}$ for all $j\in\{1,\ldots,m\}$.
Proof.
We proceed as in [Wel15] by running induction on $N$. For $N=1$, the statement is trivial. Assume the statement holds true for $N$. Consider an additional point $y_{N+1}\in\mathbb{R}^{3}$. If $\operatorname{dist}(B_{R}(y_{N+1}),B_{j})\geq R$ for all $j\in\{1,\ldots,m\}$, define $B_{m+1}:=B_{R}(y_{N+1})$ and $C_{m+1}:=\{N+1\}$.
Given $Y_{0}=(y_{1},\ldots,y_{N})$ above, we now define
@@
Notice that
@@
In particular, $\Phi_{0}$ is not totally antisymmetric, but antisymmetric only within clusters. We have proved the following result.
2.2 Lemma.
  1. (i)
    $B_{j}$ has radius $R_{j}=\frac{1}{2}(3n_{j}-1)R\quad\forall 1\leq j\leq m$,
  2. (ii)
    $\mathrm{dist}(B_{i},B_{j})\geq R$ for $i\neq j$,
  3. (iii)
2.3 Remark.
In contrast to [Wel15], the growth of the error in the lower bound now is of order $N^{2}$ and not $N$. This is due to the fact that we are accounting for fermionic statistics.
2.4 Remark.
@@
Moreover, the leading order term will be given by $E_{PT}^{(N,\alpha)}=\alpha^{2}E_{PT}^{(N,1)}$. Thus the correction from localization is, indeed, of lower order. Observe that the order of magnitude of the error in $N$ coincides with the error in Theorem 1.1.
In our next step, we will bound the total many-particle ground state energy from below by the sum of the ground state energies for less particles supported on the balls described above and the interaction between the balls. For $n\in\mathbb{N}$ and a Borel measurable set $M\subseteq\mathbb{R}^{3}$, denote by
@@
the ground state energy of an $n$-polaron whose electrons are localized in $M$.
The following proposition generalizes [FLST11, Lemma 3]. As in [Wel15], the treatment of the bipolaron is transfered to the multipolaron case. In the proof, which can be found in Appendix B, we take the fermionic statistics into account.
2.5 Proposition.
Let $\Phi_{0}$ be as constructed in Lemma 2.2 and $C_{j}$ denote the set of electrons supported in $B_{j}$. Then $\Phi_{0}$ satisfies
@@
2.6 Remark.
@@
3 UV cut-off and partition into block modes
Our goal is next to justify that one can, in fact, integrate out the phonon modes to reduce the Fröhlich model to the Pekar-Tomasevich model. For that, we follow the arguments in [LT97] very closely, by adapting the result to include external fields and interactions among the electrons, similarly to [GW13, Wel15].
@@ 3 UV cut-off and partition into block modes
$\mathcal{Q}_{n}(B)\,:=\,\{\Psi\in\Big(\bigwedge_{\ell=1}^{n}C_{c}^{\infty}(\mathbb{R}^{3})\Big)\otimes\mathcal{F}_{0}\text{ with }\operatorname{supp}(\Psi)\subset B^{n}\}.$ (3.5)
The following result is straight-forward extension of the result in [LT97] to multiple interaction electrons, see [Wel15, Lemma 4.1].
@@
For any values of $n\in\mathbb{N}$ and $\Lambda>0$ with with $4n\alpha<\pi\Lambda$ in the sense of quadratic forms on $\mathcal{Q}_{n}(B)$, we have that
@@
In every $Q(\ell)$, an arbitrary $k_{\ell}$ is chosen. The choices will be specified later. The block modes are defined by
@@
where $\mathcal{N}_{\mathrm{block}}:=\sum_{\ell\in\Lambda_{P}}a^{\dagger}_{\ell}a_{\ell}$.
Cutting into finite many blocks allows us to repeat Pekar’s argument up to a small error, as outlined after Lemma A.1 in the appendix. The details are carried out in the proof of Proposition 3.3 in Appendix C.
@@
The proof, which is presented in Appendix B below, follows analogous steps to those in [LT97], with slight adaption due to the presence of external fields, see also [Wel15].
@@
The proof essentially follows Lieb-Thomas’ argument, but we include the localized electronic support. This does not interfere with the arguments, as we only integrate phonon degrees of freedom. For more details, we refer to Appendix C.
3.4 Remark.
One of the reasons that allows us to remove Assumption (1.24) is that Proposition 3.3 is an improvement of [Wel15, Corollary 4.3]. In fact, on the right hand side of (3.11), we have the quantity $E_{PT}^{(n,\alpha,\beta^{-1})}(B)$ instead of $E_{PT}^{(n,\alpha,\beta^{-1})}(\mathbb{R}^{3})$. Together with the fact that the balls $B_{j}$ are separated with distance $R=N^{-\frac{11}{30}}\alpha^{-19/23}$, we are able to prove asymptotic subadditivity, see Lemma 3.6.
3.5 Remark.
Observe that Proposition 3.3 holds for any value $\nu\in\mathbb{R}$. This is consistent with the fact that the assumption $\nu>0$ is not really used in the proof of [Wel15, Prop. 4.2].
Notice that thus far, we kept track of the electronic support of the approximate minimizer. The electronic support is contained in a cluster of balls of comparable radii that are separated by a distance of the same order. Combined with the following Lemma, this justifies that we can remove the assumption (1.24).
3.6 Lemma.

### v3

@@ 1.3 Main result
In particular, we have that
$\lim_{\alpha\to\infty}\frac{E^{(N,\alpha)}}{\alpha^{2}}\,=\,E_{PT}^{(N,1)}\,.$ (1.27)
@@
see Lemma C.1. The infimum in (1.29) is finite, since the electron-electron interaction grows at most with $N^{2}$, while the rest grows to linear order. To bound the terms, one employs Hardy’s inequality together with the diamagnetic inequality [27, Theorem 7.21], and the assumptions on $A$ and $V$.
@@
In the absence of external fields and in the physically relevant regime $\nu>2$, Griesemer and Møller [18] established stability of the Pekar-Tomasevich functional in the sense that
@@
see also [15]. This bound is used only in the auxiliary Lemma C.1, see also (Block-mode reduction)., and improves the quadratic growth, see Remark 1.3. As a consequence, one can improve the bound the $N$-dependence of the error in Theorem 1.2 to
@@
for some constant $C_{\alpha}$ independent of $N$, see [15, Theorem 3]. In the regime $\nu<2$, this stability bound ceases to be true. In fact, by [18], one has that
@@
As a consequence of Theorem 1.2, one can establish binding of the multipolaron in the case that $V=0$ and $A$ is linear, corresponding to a constant magnetic field, as long as $\nu<\nu_{c}(A,N)$ for some threshold $\nu_{c}(A,N)>2$. More precisely, one of the authors and Griesemer [2] proved the existence of $\nu_{c}(A,N)>2$ s.t. for all $\nu<\nu_{c}(A,N)$
@@
Lewin [25] proved an analogous result in the case $A=0$. As a consequence of this result and Theorem 1.2, one obtains for all $\alpha>0$ sufficiently large
@@
As in [19], one can prove that without rescaling $A\to A_{\alpha}$, $V\to V_{\alpha}$, $\lim_{\alpha\to\infty}\alpha^{-2}E^{(N,\alpha)}$ is independent of the external fields, under sufficient regularity of $A$ and $V$.
We close the introduction with a brief proof sketch. In Section 2 we develop the cluster localization framework; this isolates the fermionic clusters and prepares the decoupling of phonon modes. Section 3 carries out the strong-coupling analysis for each cluster and derives the effective Pekar–Tomasevich energy. Section 4 assembles the estimates to prove the main theorem. The technical appendices collect the derivation of the Pekar–Tomasevich functional and the localization and phonon-integration steps.
2 Cluster localization
We begin by setting up the localization framework that underpins the remainder of the proof. Our goal is to localize an approximate minimizer of the Fröhlich Hamiltonian in a union of balls that are far from each other. The constructed minimizer will be antisymmetric within clusters contained in distinct balls, but not among different balls. Due to the spatial separation of the balls, this will allow us to decompose the Hamiltonian into independent clusters.
As usual, we denote by $B_{r}(x)\subseteq\mathbb{R}^{3}$ the open ball of radius $r$ centered at $x$.
2.1 Lemma (Covering lemma,[42]).
  1. 1.
    $B_{j}$ has radius $R_{j}=\frac{1}{2}(3n_{j}-1)R\quad\forall 1\leq j\leq m$,
  2. 2.
    $\operatorname{dist}(B_{i},B_{j})\geq R$ for $i\neq j$,
  3. 3.
    $\bigcup_{k\in C_{j}}B_{R}(y_{k})\subseteq B_{j}$ for all $j\in\{1,\ldots,m\}$.
For convenience, we included the proof of Lemma 2.1 in Appendix B.
In the following, we identify $\Psi\in\mathcal{H}_{N}$ with a vector $\iota(\Psi)\in L^{2}(\mathbb{R}^{3N};\mathcal{F})$ and denote by $\operatorname{supp}_{e}(\Psi):=\operatorname{supp}(\iota(\Psi))$ its electronic support.
2.2 Proposition (Cluster localization).
  1. 1.
    $B_{j}$ has radius $R_{j}=\frac{1}{2}(3n_{j}-1)R\quad\forall 1\leq j\leq m$,
  2. 2.
    $\mathrm{dist}(B_{i},B_{j})\geq R$ for $i\neq j$,
  3. 3.
Proof.
As explained in the introduction, we repeat the localization procedure taken from [28]. Given $\chi$ a non-negative, $L^{2}$-normalized test function localized in $B_{R}(0)$, define for all $X=(x_{1},\ldots,x_{N})$, $Y=(y_{1},\ldots,y_{N})$
@@ 2 Cluster localization
Let $\Phi\in\mathcal{H}_{N}$ be normalized. With steps analogous to those in [28, Lemma 4.1], one can show that $W$ is well-defined and smooth. Employing the IMS localization formula [11, 22, 32, 40, 41], and by approximating the Dirichlet ground-state, one then argues that there exists some $Y_{0}\in\mathbb{R}^{3N}$ such that $W_{Y_{0}}\Phi\neq 0$, and
@@
Using the Covering Lemma 2.1, we partition the electrons into clusters of possibly larger balls that are well-separated. Now define the joint wave-function of the separated electron clusters
@@
This choice of $\Phi_{0}$ satisfies the needed assumptions, completing the proof. ∎
2.3 Remark (Partial statistics).
Notice that in the proof of Prop. 2.2
@@
In particular, $\Phi_{0}$ is not totally antisymmetric, but antisymmetric only within clusters.
2.4 Remark (Error growth with $N$).
In contrast to [42], the growth of the error in the lower bound now is of order $N^{2}$ and not $N$. This is due to the fact that we are accounting for fermionic statistics.
@@
Recall from Rem. 1.1 that $E_{PT}^{(N,\alpha)}=\alpha^{2}E_{PT}^{(N,1)}$. Thus the correction from localization is, indeed, of lower order. Observe that the order of magnitude of the error in NN coincides with the error in Theorem 1.2.
In our next step, we will bound the total many-particle ground-state energy from below by the sum of the ground-state energies for less particles supported on the balls described above and the interaction between the balls.
For $n\in\mathbb{N}$ and a Borel measurable set $M\subseteq\mathbb{R}^{3}$, denote by
@@
the ground-state energy of an $n$-polaron whose electrons are localized in $M$.
The following proposition generalizes [15, Lemma 3]. As in [42], the treatment of the bipolaron is transferred to the multipolaron case. In the proof, which can be found in Appendix B, we take the fermionic statistics into account.
2.5 Proposition (Separation of clusters).
Let $\Phi_{0}$ be as constructed in Prop. 2.2 and $C_{j}$ denote the set of electrons supported in $B_{j}$. Then $\Phi_{0}$ satisfies
@@
2.6 Remark (Negligibility of inter-cluster interaction).
@@
3 Strong-coupling analysis of multipolaron clusters
With the cluster decomposition in place, we now carry out the strong-coupling analysis for each cluster and connect the Fröhlich energy to the Pekar–Tomasevich functional.
3.1 Pekar-Tomasevich reduction of individual clusters
Our next goal is to justify that one can, in fact, integrate out the phonon modes to reduce the Fröhlich model to the Pekar-Tomasevich model. For that, we follow the arguments in [29] closely, by adapting the result to include external fields and interactions among the electrons, similarly to [19, 42].
@@ 3.1 Pekar-Tomasevich reduction of individual clusters
$\mathcal{Q}_{n}(B)\,:=\,\{\Psi\in\Big(\bigwedge_{\ell=1}^{n}C_{c}^{\infty}(\mathbb{R}^{3})\Big)\otimes\mathcal{F}_{0}\text{ with }\operatorname{supp}_{e}(\Psi)\subset B^{n}\}.$ (3.5)
The following result is a straightforward extension of the result in [29] to multiple interacting electrons, see [42, Lemma 4.1].
@@
For any values of $n\in\mathbb{N}$ and $\Lambda>0$ with $4n\alpha<\pi\Lambda$ in the sense of quadratic forms on $\mathcal{Q}_{n}(B)$, we have that
@@
In every $Q(\ell)$, an arbitrary $k_{\ell}$ is chosen, and later optimized. The block modes are defined by
@@
where $\mathcal{N}_{\mathrm{block}}:=\sum_{\ell\in\Lambda_{P}}a^{\dagger}_{\ell}a_{\ell}$. Cutting into finite many blocks allows us to repeat Pekar’s argument up to a small error, as outlined after Lemma A.1 in the appendix. The details are carried out in the proof of Proposition 3.3 in Appendix C.
@@
The proof, which is presented in Appendix B below, follows analogous steps to those in [29], with slight adaption due to the presence of external fields, see also [42].
Next, we show that the ground-state energy of the Block mode Hamiltonian can be bounded below by the Pekar-Tomasevich energy, up to an error, analogous to [29]. The crucial improvement in the statement however is that we exploit the fact that the electronic support can be preserved in this bound. This insight, together with the fact that the regions, in which the clusters are localized, are well-separated, subsequently allows us to remove Assumption (1.1), see Corollary 3.7.
@@
For convenience, we carried out the proof in Appendix C. Preserving the electronic support does not interfere with the arguments in [29], as we only integrate phonon degrees of freedom. Notice that $H_{\mathrm{block}}^{(n,\alpha,\beta)}$ does not contain the entire kinetic energy, as we sacrificed a fraction of it for the UV truncation. To effectively factor out this relative factor, which converges to 1 in the strong-coupling limit, one employs a concavity argument, as previously noted in [3, 19, 42]. Since the argument is not explicitly presented in these works, we outline the complete proof in Lemma C.1 for the convenience of the reader.
3.4 Remark (Regime for effective Coulomb coupling).
Observe that Proposition 3.3 holds for any value $\nu\in\mathbb{R}$. This is consistent with the fact that the assumption $\nu>0$ is not really used in the proof of [42, Prop. 4.2].
3.2 Cluster synthesis
Notice that thus far, we have kept track of the electronic support of the approximate minimizer. The electronic support is contained in a cluster of balls of comparable radii that are separated by a distance of the same order. We will now explain how this allows us to remove Assumption (1.1). We start by explaining how the energy of disjoint clusters splits.
3.5 Lemma (Energy of disjoint clusters).
Assume that $\rho_{\psi_{1}}$, $\rho_{\psi_{2}}$, …, $\rho_{\psi_{m}}$ have disjoint support, and assume that $\psi_{j}$ is an $n_{j}$-electron wave function. Then we have that
$\hskip-4.30554pt\mathcal{E}_{PT}^{(N,\alpha)}(\psi_{1}\wedge\ldots\wedge\psi_{m})=\sum_{j=1}^{m}\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})+(\nu-2)\alpha\hskip-2.15277pt\sum_{i<j}\int\hskip-2.15277pt\int\hskip-2.15277pt\frac{\rho_{\psi_{i}}(x)\rho_{\psi_{j}}(y)}{|x-y|}\,\mathrm{d}x\,\mathrm{d}y.$ (3.13)
Proof.
Using the disjoint support of $\rho_{\psi_{i}}$ and $\rho_{\psi_{j}}$, $i\neq j$, we have that
$\rho_{\psi_{1}\wedge\ldots\wedge\psi_{m}}\,=\,\sum_{j=1}^{m}\rho_{\psi_{j}}\,.$ (3.14)
Observe that $\operatorname{supp}(\psi_{j})\subseteq\operatorname{supp}(\rho_{\psi_{j}})^{n_{j}}$. For every local operator $\mathcal{O}$ that is symmetric with respect to the coordinates $x_{1},\dots,x_{N}$ we have that
$\langle\psi_{1}\wedge\ldots\wedge\psi_{m},\mathcal{O}\psi_{1}\wedge\ldots\wedge\psi_{m}\rangle=\langle\psi_{1}\otimes\ldots\otimes\psi_{m},\mathcal{O}\psi_{1}\otimes\ldots\otimes\psi_{m}\rangle\,.$ (3.15)
As a consequence, decomposing the interaction terms to the ones that are internal within a cluster and interaction terms between different clusters, we obtain
$\mathcal{E}_{PT}^{(N,\alpha)}(\psi_{1}\wedge\ldots\wedge\psi_{m})\,$ $=\,\sum_{j=1}^{m}\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})-\alpha\sum_{i\neq j}\int_{B_{i}}\int_{B_{j}}\frac{\rho_{\psi_{i}}(x)\rho_{\psi_{j}}(y)}{|x-y|}\,\mathrm{d}x\,\mathrm{d}y$ (3.16)
where $C_{i},C_{j}$ denote the $i$-th and $j$-th clusters, and the factor $\frac{1}{2}$ stems from replacing $\sum_{i<j}$ by $\sum_{i\neq j}$. Using the fact that $\rho_{\psi_{i}}(x)=|C_{i}|\int_{\mathbb{R}^{3(|C_{i}|-1)}}|\psi_{i}(x,X)|^{2}\,\mathrm{d}X$, we arrive at (3.13). ∎
3.6 Remark (Translation-invariance and subadditivity of $E_{PT}$).
To understand the subtlety of our new auxiliary result Corollary 3.7, let us explain the argument to prove (1.1) in the external field-free case. Let $\psi_{j}$ be an approximate minimizer of the Pekar-Tomasevich functional $E_{PT}^{(n_{j},\alpha)}$ with compact support, up to error $\frac{\varepsilon}{m}$. Defining the translation operator $(\tau_{z}f)(x_{1},\ldots,x_{\ell}):=f(x_{1}+z,\ldots,x_{\ell}+z)$, we consider the sequence of test functions
$\Psi_{k}\,:=\,\tau_{k\hat{e}_{1}}\psi_{1}\wedge\tau_{k^{2}\hat{e}_{1}}\psi_{2}\wedge\ldots\wedge\tau_{k^{m}\hat{e}_{1}}\psi_{m}\,.$ (3.18)
For $k$ sufficiently large, we have that the supports of $\tau_{k^{i}\hat{e}_{1}}\psi_{i}$, $\tau_{k^{j}\hat{e}_{1}}\psi_{j}$, $i\neq j$, are disjoint. Using Lemma 3.5, we then obtain
$E_{PT}^{(N,\alpha)}\,\leq\,\liminf_{k\to\infty}\mathcal{E}_{PT}^{(N,\alpha)}(\Psi_{k})\,\leq\,\sum_{j=1}^{m}\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})\,\leq\,\sum_{j=1}^{m}E_{PT}^{(n_{j},\alpha)}+\varepsilon\,.$ (3.19)
Since external fields break translation-invariance of $E_{PT}^{(n_{j},\alpha)}$, we now show how to adapt the argument to our present setting. To accomplish that, we will exploit the fact that the localized clusters are well-separated.
3.7 Corollary (Asymptotic subadditivity of $E_{PT}$).